OpenStudy (anonymous):

Simplify the fraction a^3 + b^3/(a+b)^3

OpenStudy (ajprincess):

\((a^3+b^3)=(a+b)(a^2-ab+b^2)\) \((a+b)^3=(a+b)(a^2+2ab+b^2)\) Cancel out the common factors and u wil get the simplified form. Hope that helps:)

OpenStudy (anonymous):

That means I would have to cancel out a+b

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

Thanks

OpenStudy (ajprincess):

welcome:)